OpenStudy (anonymous):

Find y''(x) by implicit differentiation from sqrt(x)+sqrt(y)=1

OpenStudy (anonymous):

\[ \frac{y'(x)}{2 \sqrt{y(x)}}+\frac{1}{2 \sqrt{x}}=0\\ y'(x)=-\frac{\sqrt{y(x)}}{\sqrt{x}} \]

OpenStudy (anonymous):

Can you find y''(x)?

OpenStudy (anonymous):

i can get for y'....but i stuck when to find y''....

OpenStudy (cggurumanjunath):